OpenStudy (anonymous):

Find the solutions of the inequality. | x - 5 | ≥ 3

OpenStudy (anonymous):

since absolute value has two solutions one positive and one negative. \[\Large x-5\ge3\] and \[\Large x-5\ge-3\] solving the first \[\Large x \ge3+5\] \[\Large x \ge8\] can you solve the second equation ?? \[\Large x-5\ge-3\] ?

OpenStudy (anonymous):

it looks strange i cnt see it. is says large x-5

OpenStudy (anonymous):

@Pavlyunchenko

OpenStudy (anonymous):

the whole thing change

OpenStudy (anonymous):

are you ok with the first equation solution?

OpenStudy (anonymous):

it wont appear as before. its like text change

OpenStudy (anonymous):

ge-3? what is that

OpenStudy (anonymous):

it is necessary to write as ge-3 because we have removed the absolute signs and absolute signs have two have values jo it should be ge 3 and ge-3

hero (hero):

@sami-21, your approach is not correct

hero (hero):

You're too confident in your approach

OpenStudy (anonymous):

yes i am sure .

hero (hero):

| x - 5 | ≥ 3 becomes: x - 5 ≥ 3 or x - 5 ≤ -3

OpenStudy (anonymous):

no

hero (hero):

Yes

hero (hero):

I assume you think you know what the solution is

hero (hero):

That's even more incorrect

hero (hero):

You have the inequality signs facing the wrong way

hero (hero):

If you use what I suggested, you get \(x \le 2\) \(x \ge 8\) as solutions which wolframalpha confirms: http://www.wolframalpha.com/input/?i=solve+%28%7C+x+-+5+%7C+%E2%89%A5+3%2Cx%29

OpenStudy (anonymous):

sorry ! yes you are right

hero (hero):

I arrived at my equations by using the simple rule that \(|x-a| \ge b\) is equivalent to \(-b \ge x - a \ge b\), then separate into two equations from there: \(x - a \ge b\) \(x-a \le - b\)

OpenStudy (anonymous):

yes thats corret !

OpenStudy (anonymous):

thansk so is both the answers x - 5 ≥ 3 or x - 5 ≤ -3

hero (hero):

You need to isolate x in order to find the solutions

OpenStudy (anonymous):

srry yea is x ≥ 8 right?

hero (hero):

That's one solution